# POLAR: A Pessimistic Model-based Policy Learning Algorithm for Dynamic Treatment Regimes

### v1

@@ C.1.2 Proof of Proposition 1
where $\widetilde{O}$ hides all logarithmic terms and $(i)$ comes from Corollary S1.
@@
[…]
@@
Next, by Lemma S11, we have $\mathbb{E}_{x\sim\rho}\widehat{h}(x,x)\lesssim n^{\frac{1}{\omega}-1},$ which implies
@@ D.1 Supporting Lemmas and Their Proofs of Proposition S1
Thus, $X$ is a zero-mean, bounded random variable supported on $[-\sigma,\sigma]$. We now verify that $X$ is $\sigma$-sub-Gaussian. By Definition S1(i), we must show that for all $\lambda\in\mathbb{R}$, $\mathbb{E}[e^{\lambda X}]\leq\exp\left(\frac{\lambda^{2}\sigma^{2}}{2}\right).$ Since $\lambda X\in[-|\lambda|\sigma,|\lambda|\sigma]$, we apply the convexity of the exponential function:
@@ D.3 Supporting Lemmas’ Proofs of Theorem S2
=$\stackrel{{\scriptstyle(i)}}{{=}}\left\langle Q_{t}^{\widehat{\pi}}(h_{t},\cdot),\ \pi_{t}(\cdot\mid h_{t})-\widehat{\pi}_{t}(\cdot\mid h_{t})\right\rangle,$
@@
$-\operatorname{KL}(\pi_{k}^{(t)},\pi_{k}^{(t+1)})-\operatorname{KL}(\pi_{k}^{(t+1)},\pi_{k}^{(t)})=\left\langle f_{k}(\theta_{k}^{(t+1)},h_{k},\cdot)-f_{k}(\theta_{k}^{(t)},h_{k},\cdot),\ \pi_{k}^{(t)}-\pi_{k}^{(t+1)}\right\rangle,$ (S46)
@@ D.4 Supporting Lemmas and Their Proofs for Theorem S3
Since $\mathcal{X}$ is compact and $f\in C^{1}(\mathcal{X})$, we define $C_{i}:=\sup_{x\in\mathcal{X}}|f_{i}^{\prime}(x)|<\infty$, and let $C:=\sum_{i=1}^{n}C_{i}$. Then for any $x,y\in\mathcal{X}$, write $x-y=sv$ for some $s=\|x-y\|_{2}\geq 0$ and unit vector $v$. Since $f\in C^{1}$,
[…]
@@
[…]
@@
Suppose Assumptions S3, S4, S5 hold, then $Q_{k,\widetilde{M}}^{\pi^{(t)}}(h_{k},a_{k})$ is $p$-smooth in $(s_{1},s_{2},...,s_{k})$ for any $10\leq t\leq T-$ 1, $1\leq k\leq K$, and any fixed $(a_{1},...,a_{k})\in\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{k}$.
@@
Recall that the Algorithm 2 only update the policy within the policy class $\Pi$, so $\pi^{(t)}\in\Pi$ for any $10\leq t\leq T-$ 1. We only need to prove $Q_{k,\widetilde{M}}^{\pi}(h_{k},a_{k})$ is $p$-smooth in $(s_{1},s_{2},...,s_{k})$ for any $\pi\in\Pi$, any $k$, any $(a_{1},...,a_{k})\in\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{k}$.
@@
Now we focus on the induction step from $k+1$ to $k$. Suppose $Q_{k+1,\widetilde{M}}^{\pi}(h_{k+1},a_{k+1})$ is $p$-smooth in $(s_{1},...,s_{k+1})$ for some $11\leq k\leq K-$ 1. By Bellman equation,
[…]
@@ D.5 Supporting Lemmas’ Proofs for Theorem 3
[…]
@@
[…]
@@
[…]
@@ Appendix E Simulation Setup
The noises are defined as $\varepsilon_{k}^{i}=0.8\left(z_{k}^{i}-0.5\right),\text{ where }z_{k}^{i}\sim\operatorname{Beta}(2,2)\text{ so that }\varepsilon_{k}^{i}\in[-0.4,0.4]$, $i=1,2$.

### v2

@@ C.1.2 Proof of Proposition 1
where $\widetilde{O}$ hides all logarithmic terms and $(i)$ comes from Corollary S1. ∎
Proof of Theorem 3.
Define the index set $\mathcal{U}:=\prod_{k=1}^{K}\{-1,+1\}^{k}=\big\{u=(u_{1},\dots,u_{K}):u_{k}\in\{-1,+1\}^{k}\big\},$ and let $m:=\sum_{k=1}^{K}k=\frac{K(K+1)}{2}.$ For each $u\in\mathcal{U}$ we construct a DTR model $M_{u}$ with linear transition kernels as follows. The transition kernels follow the linear structure and do not depend on $u$:
$\mathbb{E}[S_{k+1}\mid H_{k},A_{k}]\;=\;W_{k}\,\phi_{k}(H_{k},A_{k}),\qquad k=1,\dots,K,$
for some weight matrices $W_{k}$ and feature mapping $\phi_{k}:\mathcal{H}_{k}\times\mathcal{A}_{k}\to\mathbb{R}^{\dim(\phi_{k})}$ satisfying $\dim(\phi_{k})\lesssim k$. In our hard instance, we choose $\dim(\phi_{k})=k$ and, at each stage $k$, set the action space to $\mathcal{A}_{k}=\{-1,+1\}^{k}$ and define $\phi_{k}(H_{k},a_{k}):=a_{k}\in\mathbb{R}^{k},\;\;a_{k}\in\mathcal{A}_{k}.$
The behavior policy is chosen to be uniform over actions: $\mu_{k}(a_{k}\mid H_{k})=2^{-k},\;\;a_{k}\in\mathcal{A}_{k},$ for all stages $k$ and all histories $H_{k}$. The mean reward at stage $k$ is assumed to be linear in the same feature mapping $\phi_{k}$:
$r_{u,k}(H_{k},A_{k})\;=\;r_{0,k}(H_{k},A_{k})\;+\;\frac{\zeta}{\sqrt{m}}\,\big\langle\phi_{k}(H_{k},A_{k}),u_{k}\big\rangle\;=\;r_{0,k}(H_{k},A_{k})\;+\;\frac{\zeta}{\sqrt{m}}\,\big\langle A_{k},u_{k}\big\rangle,$
for a scalar $\zeta>0$ to be specified later. The observed reward is $R_{u,k}(H_{k},A_{k})\sim\mathcal{N}\big(r_{u,k}(H_{k},A_{k}),\,1\big),$ independently across stages and trajectories. We denote by $\mathcal{M}:=\{M_{u}:u\in\mathcal{U}\}.$
For any deterministic policy $\pi$, we associate a sign pattern $u^{\pi}=(u_{1}^{\pi},\dots,u_{K}^{\pi})$ with $u_{k}^{\pi}\in\{-1,+1\}^{k}$ as follows: at stage $k$, $\pi$ chooses some action $A_{k}^{\pi}(H_{k})\in\{-1,+1\}^{k}$, and we set $u_{k}^{\pi}:=A_{k}^{\pi}(H_{k}).$ We then define the Hamming distance as follows
$D_{H}(u^{\pi},u):=\sum_{k=1}^{K}\sum_{j=1}^{k}\mathbf{1}\!\big\{u_{k}^{\pi}[j]\neq u_{k}[j]\big\},\qquad u\in\mathcal{U}.$
For a fixed $u$ and policy $\pi$, let $V_{M_{u}}^{\pi}$ denote its undiscounted expected cumulative value. At stage $k$, given the history $H_{k}$, the optimal policy $\pi^{*}$ for $M_{u}$ chooses the action $A_{k}^{\pi^{*}}(H_{k})=u_{k},$ which maximizes $\langle A_{k},u_{k}\rangle$ over $A_{k}\in\{-1,+1\}^{k}$. In contrast, policy $\pi$ chooses $A_{k}^{\pi}(H_{k})=u_{k}^{\pi}$. Since $r_{0,k}$ does not depend on $u$, we have the reward difference:
$r_{u,k}(H_{k},A_{k}^{\pi^{*}}(H_{k}))-r_{u,k}(H_{k},A_{k}^{\pi}(H_{k}))=\frac{\zeta}{\sqrt{m}}\Big(\langle u_{k},u_{k}\rangle-\langle u_{k},u_{k}^{\pi}\rangle\Big).$
Using $u_{k}[j],u_{k}^{\pi}[j]\in\{-1,+1\}$, we compute
$\langle u_{k},u_{k}\rangle-\langle u_{k},u_{k}^{\pi}\rangle=\sum_{j=1}^{k}\big(u_{k}[j]^{2}-u_{k}[j]u_{k}^{\pi}[j]\big)=\sum_{j=1}^{k}\big(1-u_{k}[j]u_{k}^{\pi}[j]\big)=2\sum_{j=1}^{k}\mathbf{1}\{u_{k}[j]\neq u_{k}^{\pi}[j]\}.$
Hence, for every history $H_{k}$,
$r_{u,k}(H_{k},A_{k}^{\pi^{*}}(H_{k}))-r_{u,k}(H_{k},A_{k}^{\pi}(H_{k}))\;=\;\frac{2\zeta}{\sqrt{m}}\sum_{j=1}^{k}\mathbf{1}\{u_{k}^{\pi}[j]\neq u_{k}[j]\}.$
Taking expectation and summing over $k=1,\dots,K$, we obtain that
$V_{M_{u}}^{\pi^{*}}-V_{M_{u}}^{\pi}=\sum_{k=1}^{K}\mathbb{E}_{M_{u}}\big[r_{u,k}(H_{k},A_{k}^{\pi}(H_{k}))-r_{u,k}(H_{k},A_{k}^{\pi}(H_{k}))\big]=\frac{2\zeta}{\sqrt{m}}\sum_{k=1}^{K}\sum_{j=1}^{k}\mathbf{1}\{u_{k}^{\pi}[j]\neq u_{k}[j]\}.$
Recall the definition of $D_{H}(u^{\pi},u)$, this can be rewritten as
$V_{M_{u}}^{\pi^{*}}-V_{M_{u}}^{\pi}\;\geq\;\gamma_{\mathrm{v}}\,\frac{\zeta}{\sqrt{m}}\,D_{H}(u^{\pi},u),$ (S33)
where we may take $\gamma_{\mathrm{v}}:=2$.
Now we bound the KL divergence between neighboring models. Let $Q_{u}$ denote the distribution of $\mathcal{D}$ under $M_{u}$, and let $P_{u}^{(1)}$ be the distribution of a single trajectory under $M_{u}$. If $u,u^{\prime}\in\mathcal{U}$ satisfy $D_{H}(u,u^{\prime})=1$, then the corresponding models $M_{u}$ and $M_{u^{\prime}}$ have identical transition kernels, but their reward parameters differ in exactly one component, say at stage $k^{\star}$ and coordinate $\ell^{\star}$ of $u_{k^{\star}}$.
At stage $k^{\star}$, the reward distributions under $M_{u}$ and $M_{u^{\prime}}$ conditioned on the action $A_{k^{\star}}=a\in\{-1,+1\}^{k}$ are
$R_{u,k^{\star}}\mid A_{k^{\star}}=a\sim\mathcal{N}\!\left(r_{0,k^{\star}}(H_{k^{\star}},a)+\frac{\zeta}{\sqrt{m}}\langle u_{k^{\star}},a\rangle,\;1\right),$
$R_{u^{\prime},k^{\star}}\mid A_{k^{\star}}=a\sim\mathcal{N}\!\left(r_{0,k^{\star}}(H_{k^{\star}},a)+\frac{\zeta}{\sqrt{m}}\langle u^{\prime}_{k^{\star}},a\rangle,\;1\right),$
while at all other stages $k\neq k^{\star}$ the reward distributions coincide. Since $u_{k^{\star}}$ and $u^{\prime}_{k^{\star}}$ differ only in coordinate $\ell^{\star}$, we have $u_{k^{\star}}-u^{\prime}_{k^{\star}}=\pm 2e_{\ell^{\star}}.$ Therefore, for any $a\in\{-1,+1\}^{k}$,
$\langle u_{k^{\star}},a\rangle-\langle u^{\prime}_{k^{\star}},a\rangle=(u_{k^{\star}}[\ell^{\star}]-u^{\prime}_{k^{\star}}[\ell^{\star}])\,a[\ell^{\star}]\in\{\pm 2\}.$
Therefore, conditioned on $A_{k^{\star}}=a$,
$\mathrm{KL}\!\Big(\text{Law}(R_{u,k^{\star}}\mid A_{k^{\star}}=a)\,\big\|\,\text{Law}(R_{u^{\prime},k^{\star}}\mid A_{k^{\star}}=a)\Big)=\frac{1}{2}\cdot\frac{4\zeta^{2}}{m}=\frac{2\zeta^{2}}{m},$
and this quantity does not depend on $a$. Taking expectation over the random action $A_{k^{\star}}$ yields
$\mathrm{KL}\big(P_{u}^{(1)}\,\|\,P_{u^{\prime}}^{(1)}\big)=\frac{2\zeta^{2}}{m}.$
Because the $n$ trajectories in $\mathcal{D}$ are i.i.d., the KL divergence between $Q_{u}$ and $Q_{u^{\prime}}$ is
$\mathrm{KL}(Q_{u}\,\|\,Q_{u^{\prime}})=n\,\mathrm{KL}\big(P_{u}^{(1)}\,\|\,P_{u^{\prime}}^{(1)}\big)=\frac{2n\zeta^{2}}{m}.$ (S34)
Now choose $\zeta^{2}=\frac{\alpha_{\mathrm{KL}}m}{n}$ for some sufficiently small constant $\alpha_{\mathrm{KL}}>0$. Then (S34) gives
$\mathrm{KL}(Q_{u}\,\|\,Q_{u^{\prime}})\;<\;2\alpha_{\mathrm{KL}}=:C_{\mathrm{KL}},$
for all neighbors $u,u^{\prime}$ with $D_{H}(u,u^{\prime})=1$, where $C_{\mathrm{KL}}$ is a fixed constant. This uniform bound on the KL divergences of neighboring models is exactly the condition required in Assouad’s lemma (Tsybakov,, 2009).
By Assouad’s lemma (Lemma 2.12 in Tsybakov, (2009)), there exists a constant $\alpha_{\mathrm{A}}>0$ (depending only on $C_{\mathrm{KL}}$) such that, for the final policy $\widehat{\pi}$ of any learning algorithm,
$\sup_{u\in\mathcal{U}}\mathbb{E}_{u}\big[D_{H}(u^{\widehat{\pi}},u)\big]\;\geq\;\alpha_{\mathrm{A}}\,m,$ (S35)
where $m:=\sum_{k=1}^{K}k=\frac{K(K+1)}{2}.$
Since $0\leq D_{H}(u^{\widehat{\pi}},u)\leq m$ for any fixed $u\in\mathcal{U}:=\prod_{k=1}^{K}\{-1,+1\}^{k}$,
$\mathbb{E}_{u}\big[D_{H}(u^{\widehat{\pi}},u)\big]$ $=\mathbb{E}_{u}\!\left[D_{H}(u^{\widehat{\pi}},u)\,\mathbf{1}\!\left\{D_{H}(u^{\widehat{\pi}},u)<\frac{\alpha_{\mathrm{A}}}{2}m\right\}\right]$
$+\mathbb{E}_{u}\!\left[D_{H}(u^{\widehat{\pi}},u)\,\mathbf{1}\!\left\{D_{H}(u^{\widehat{\pi}},u)\geq\frac{\alpha_{\mathrm{A}}}{2}m\right\}\right]$
$\leq\frac{\alpha_{\mathrm{A}}}{2}m+m\,\mathbb{P}_{u}\!\left(D_{H}(u^{\widehat{\pi}},u)\geq\frac{\alpha_{\mathrm{A}}}{2}m\right).$
Taking the supremum over $u$ and using $\sup_{u}\mathbb{E}_{u}[D_{H}(u^{\widehat{\pi}},u)]\geq\alpha_{\mathrm{A}}m$, we obtain
$\alpha_{\mathrm{A}}m\;\leq\;\frac{\alpha_{\mathrm{A}}}{2}m+m\sup_{u\in\mathcal{U}}\mathbb{P}_{u}\!\left(D_{H}(u^{\widehat{\pi}},u)\geq\frac{\alpha_{\mathrm{A}}}{2}m\right),$
so there exists some $u^{\star}\in\mathcal{U}$ such that
$\mathbb{P}_{u^{\star}}\!\left(D_{H}(u^{\widehat{\pi}},u^{\star})\geq\frac{\alpha_{\mathrm{A}}}{2}m\right)\;\geq\;\frac{\alpha_{\mathrm{A}}}{2}.$ (S36)
Using (S33) with $\pi=\widehat{\pi}$ and $u=u^{\star}$,
$V_{M_{u^{\star}}}^{\pi^{*}}-V_{M_{u^{\star}}}^{\widehat{\pi}}\;\geq\;\gamma_{\mathrm{v}}\frac{\zeta}{\sqrt{m}}\,D_{H}(u^{\widehat{\pi}},u^{\star}),$
and on the event in (S36) this yields
$V_{M_{u^{\star}}}^{\pi^{*}}-V_{M_{u^{\star}}}^{\widehat{\pi}}\;\geq\;\gamma_{\mathrm{v}}\frac{\zeta}{\sqrt{m}}\cdot\frac{\alpha_{\mathrm{A}}}{2}m=\frac{\gamma_{\mathrm{v}}\alpha_{\mathrm{A}}}{2}\,\zeta\sqrt{m}.$
With $\zeta^{2}=\alpha_{\mathrm{KL}}m/n$ and $m\asymp K^{2}$, there exists a universal constant $C_{\text{lower}}:=\alpha_{\mathrm{A}}\,\sqrt{\alpha_{\mathrm{KL}}}>0$ such that
$\mathbb{P}_{M_{u^{\star}}}\!\left(V_{M_{u^{\star}}}^{\pi^{*}}-V_{M_{u^{\star}}}^{\widehat{\pi}}\;\geq\;C_{\text{lower}}\frac{K^{2}}{\sqrt{n}}\right)\;\geq\;\frac{\alpha_{\mathrm{A}}}{2}.$
If we pick $\alpha_{\mathrm{A}}=1/2$, this shows that for any algorithm $\mathcal{A}$,
$\sup_{M\in\mathcal{M}}\mathbb{P}_{M}\!\left(V_{M}^{\pi^{*}}-V_{M}^{\widehat{\pi}}\;\geq\;C_{\text{lower}}\frac{K^{2}}{\sqrt{n}}\right)\;\geq\;\frac{1}{4},$
and taking $\inf_{\mathcal{A}}$ over all algorithms yields the theorem. ∎
@@
$\int_{\mathcal{X}}h(x,x^{\prime})\psi_{j}(x^{\prime})\,\rho(x^{\prime})\,dx^{\prime}=\mu_{j}\psi_{j}(x),\quad\text{for all }j\geq 1.$ (S37)
@@
Next, by Lemma S11, we have $\mathbb{E}_{x\sim\rho^{k}}\widehat{h}(x,x)\lesssim n^{\frac{1}{\omega}-1},$ which implies
@@ D.1 Supporting Lemmas and Their Proofs of Proposition S1
Thus, $X$ is a zero-mean, bounded random variable supported on $\sigma,\sigma]$. We now verify that $X$ is $\sigma$-sub-Gaussian. By Definition S1(i), we must show that for all $\lambda\in\mathbb{R}$, $\mathbb{E}[e^{\lambda X}]\leq\exp\left(\frac{\lambda^{2}\sigma^{2}}{2}\right).$ Since $\lambda X\in[-|\lambda|\sigma,|\lambda|\sigma]$, we apply the convexity of the exponential function:
@@ D.3 Supporting Lemmas’ Proofs of Theorem S2
$\stackrel{{\scriptstyle(i)}}{{=}}\left\langle Q_{t}^{\widehat{\pi}}(h_{t},\cdot),\ \pi_{t}(\cdot\mid h_{t})-\widehat{\pi}_{t}(\cdot\mid h_{t})\right\rangle,$
@@
$\operatorname{KL}(\pi_{k}^{(t)},\pi_{k}^{(t+1)})-\operatorname{KL}(\pi_{k}^{(t+1)},\pi_{k}^{(t)})=\left\langle f_{k}(\theta_{k}^{(t+1)},h_{k},\cdot)-f_{k}(\theta_{k}^{(t)},h_{k},\cdot),\ \pi_{k}^{(t)}-\pi_{k}^{(t+1)}\right\rangle,$ (S50)
@@ D.4 Supporting Lemmas and Their Proofs for Theorem S3
Since $\mathcal{X}$ is compact and $f\in C^{1}(\mathcal{X})$, we define $C_{i}:=\sup_{x\in\mathcal{X}}|f_{i}^{\prime}(x)|<\infty$, and let $C:=\sum_{i=1}^{n}C_{i}$. Then for any $x,y\in\mathcal{X}$, write xy=svx-y=sv for some $s=\|x-y\|_{2}\geq 0$ and unit vector $v$. Since $f\in C^{1}$,
$f(x)-f(y)=\int_{0}^{s}f^{\prime}_{v}(y+tv)\,dt.$
@@
$|D^{\kappa}h(x)-D^{\kappa}h(y)|$ $=\left|\int_{\mathcal{X}_{n}}\left[D^{\kappa}f(x,x_{n})-D^{\kappa}f(y,x_{n})\right]dx_{n}\right|$
$\leq\int_{\mathcal{X}_{n}}\left|D^{\kappa}f(x,x_{n})-D^{\kappa}f(y,x_{n})\right|dx_{n}$
$\leq\int_{\mathcal{X}_{n}}c\|x-y\|_{2}^{\beta}dx_{n}$
@@
Suppose Assumptions S3, S4, S5 hold, then $Q_{k,\widetilde{M}}^{\pi^{(t)}}(h_{k},a_{k})$ is $p$-smooth in $(s_{1},s_{2},...,s_{k})$ for any $0\leq t\leq T-1$, $1\leq k\leq K$, and any fixed $(a_{1},...,a_{k})\in\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{k}$.
@@
Recall that the Algorithm 2 only update the policy within the policy class $\Pi$, so $\pi^{(t)}\in\Pi$ for any $0\leq t\leq T-1$. We only need to prove $Q_{k,\widetilde{M}}^{\pi}(h_{k},a_{k})$ is $p$-smooth in $(s_{1},s_{2},...,s_{k})$ for any $\pi\in\Pi$, any $k$, any $(a_{1},...,a_{k})\in\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{k}$.
@@
Now we focus on the induction step from $k+1$ to $k$. Suppose $Q_{k+1,\widetilde{M}}^{\pi}(h_{k+1},a_{k+1})$ is $p$-smooth in $(s_{1},...,s_{k+1})$ for some $1\leq k\leq K-1$. By Bellman equation,
$\begin{split}Q_{k,\widetilde{M}}^{\pi}(h_{k},a_{k})=&\ \widetilde{r}_{k}(h_{k},a_{k})\\ &+\int_{\mathcal{S}_{k}}\left(\sum_{a_{k+1}\in\mathcal{A}_{k+1}}\pi_{k+1}(a_{k+1}|h_{k+1})Q_{k+1,\widetilde{M}}^{\pi}(h_{k+1},a_{k+1})\right)\widehat{p}_{k}(s_{k+1}|h_{k},a_{k})ds_{k+1}.\end{split}$ (S56)
@@ D.5 Supporting Lemmas’ Proofs for Theorem 4
$\int_{\mathcal{X}\times\mathcal{X}}h(x,x^{\prime})\,\rho(x)\rho(x^{\prime})\,dx\,dx^{\prime}<\infty.$
@@
$\sum_{j=1}^{\infty}\frac{2\sigma^{2}\mu_{j}}{2\sigma^{2}+n\mu_{j}}\lesssim\frac{\sigma^{2}}{n}j_{0}+\sum_{j=j_{0}+1}^{\infty}\mu_{j}\lesssim n^{-1}\cdot\left(\frac{n}{\sigma^{2}}\right)^{1/\omega}+\int_{j_{0}}^{\infty}j^{-\omega}\,dj=\mathcal{O}(n^{1/\omega-1}).$
@@
Under Assumption S10, the eigenvalues satisfy $\sum_{i=J+1}^{\infty}\mu_{i}\leq\int_{J}^{\infty}i^{-\omega}\,di\lesssim J^{-(\omega-1)}$. This implies that
@@ Appendix E Simulation Setup
[…]